A 2-m by 3-m rug is spread inside a tepee that has a diameter of x meters. Find a polynomial for the area of the tepee’s floor that is not covered.
Answer by mananth(16946)

A 2-m by 3-m rug is spread inside a tepee that has a diameter of x meters. Find a polynomial for the area of the tepee’s floor that is not covered.
...
area of rug = 2*3 = 6m^2
area of teepee = pi*(x/2)^2
=pix^2/4
Balanxce area = pi*x^2/4 - 6
...
m.ananth@hotmail.ca